OpenStudy (anonymous):

Evaluate the following integral:

OpenStudy (campbell_st):

well on initial observation, the numerator looks like the derivative of the denominator.. does that help..?

OpenStudy (anonymous):



OpenStudy (anonymous):

I want to set u= x^2 + 1 I think

OpenStudy (anonymous):

then I think dx= du/2x

OpenStudy (campbell_st):

ok... so a change in question so its

OpenStudy (campbell_st):

that's correct so looking at the derivative \[\frac{1}{2} du = x~ dx\] so you are looking at \[\frac{1}{2} \int\limits \frac{1}{u^2} ~de = \frac{1}{2} \int\limits u^{-2} du\]

OpenStudy (campbell_st):

now you should be able to find the derivative

OpenStudy (anonymous):

I received 1/2 * (x^2 + 1)^(-1) / -1